# P. Guiraud, M. Hernández, A. Meyroneinc and A. Nogueira

### v1

@@ 2.2 Main results
Let $\lambda\in(0,1)$ and $d\in(0,1-\lambda)$. For any $\rho\in(0,1)$ and $\alpha\in[0,1]$, let $\delta(\rho,\alpha)$ and $a(\delta,\rho,\alpha)$ be defined by
@@
    if $k\rho\notin\mathbb{Z}+\alpha$ for every $k\in\mathbb{Z}^{*}$ and either $\delta(\rho^{-},\alpha)<\delta(\rho,\alpha)$ or $a(\delta,\rho^{+},\alpha)<a(\delta,\rho,\alpha)$, otherwise.
  2. 2.
@@
  3. 3.
    We have $[\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\subset(1-\lambda-d,1)$, and if $\delta$ satisfies (12), then $[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\cap[0,1]\neq\emptyset$ with $[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[0,1]$ if $\alpha\in(0,1)$ or $\rho\notin\mathbb{Q}$.
@@
Theorem 2.3 can be read as follows. Let $\rho\in(0,1)$. If $\alpha$ is any number in $[0,1]$ and $\delta$ is chosen in $[\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\cap(1-\lambda-d,1)$, then, to obtain a map (3) with rotation number $\rho$ it is enough to select any $a\in[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\cap[0,1]$. Items 1) and 2) ensure in particular that the region of the parameters $\delta$ and $a$ defined by (12) is nonempty for any $\rho\in(0,1)$ and $\alpha\in[0,1]$. Then, item 3) shows that this region intersects our set of interest $M=(1-\lambda-d,1)\times[0,1]$ and is actually included in $M$, unless $\rho\in\mathbb{Q}$ and $\alpha\in\{0,1\}$.
If $\rho=p/q$ is rational, the series defining $\delta(\rho,\alpha)$ and $a(\delta,\rho,\alpha)$ are finite sums given in Lemma 4.9, which can be numerically computed with a high precision. Also, $\delta(\rho^{-},\alpha)$ and $a(\delta,\rho^{+},\alpha)$ are deduced from item 2), or from Lemma 4.6 if $rp/q\in\mathbb{Z}+\alpha$ for some $r\in\{1,\dots,q-1\}$. As an example, we carry out the numerical computation of the region of parameters defined by (12) for $\rho=1/3$ when $\alpha=1/2$ and plot it in the left panel of Figure 4. Since $a(\delta,\rho,\alpha)$ is an affine function of $\delta$ and $a(\delta,\rho,\alpha)-a(\delta,\rho^{+},\alpha)$ does not depend on $\delta$ (see item 2 or Lemma 4.6), the region obtained is a parallelogram. Now, as $\delta(p/q,\alpha)$ and $a(\delta,p/q,\alpha)$ are piecewise constant functions of $\alpha$, see Lemma 4.9, if $\alpha$ takes all the values in $[0,1]$, then (12) gives only a finite number of different regions for which the rotation number is $p/q$. If $p$ and $q$ are coprime, this number is equal to $12q+$ 1 and the regions are obtained by letting $\alpha=l/q$ for every $l\in\{0,\dots,q\}$ and by choosing any $\alpha\in(l/q,(l+1)/q)$ for every $l\in\{0,\dots,q-1\}$. In the right panel of Figure 4, the $72q+1=$ 7 parallelograms corresponding to $\rho=1/3$ are shown. We observe that the parallelograms obtained for $\alpha=l/q$ and $\alpha=(l+1)/q$ have a non-empty intersection which is the parallelogram obtained for any $\alpha\in(l/q,(l+1)/q)$.
Figure 4: In green, the values of $(\delta,a)$ in the region defined by (12) for $\lambda=0.7,d=0.2$ and $\rho=1/3$. Left panel: the parallelogram corresponding to $\alpha=1/2\in(1/3,2/3)$. In the right panel $\alpha$ takes all the values in $[0,1]$. We observe that there is $72q+1=$ 7 different parallelograms and that the parallelogram of the left panel is the intersection of two of them.
@@
As shown by items 1) and 2), the size of the intervals (12) depends on the rational or irrational nature of $\rho$, but also on the relation between $\rho$ and $\alpha$. For $\rho$ irrational, the intervals (12) are both a singleton for almost all values of $\alpha$. These intervals are not singletons only if $\rho$ and $\alpha$ are such that $\{k\rho\}=\alpha$ for some $k\in\mathbb{Z}^{*}$, that is, if $\alpha$ belongs to the orbit of $\rho$ by the irrational rotation of the same angle. If for some $\alpha$ such a $k_{\alpha}$ exits, then it is unique and one of the two intervals (but not both) is not a singleton. Its exact length depends on $k_{\alpha}$ and is given in Lemma 4.5. For $\rho=p/q$ rational, the intervals (12) have always positive Lebesgue measure. The inequalities (13) and (14) become equalities if $\alpha$ does not belong to the $q$-periodic orbit of $\rho$ by the rotation of the same angle. In particular, if $p$ and $q$ are coprime, the size of both intervals is constant in $[0,1]\setminus\{0,1/q,\dots,1\}$. Now, for $\alpha\in\{0,1/q,\dots,1\}$ the intervals are bigger and their exact length, which depends on $\alpha$, is given in Lemma 4.6. This explains the relative sizes of the parallelograms of Figure 4.
Remark 2.4.
We note that in the two-dimensional subspace of parameters where $\lambda$ and $d$ are fixed, the regions of $(\delta,a)$ corresponding to rational rotation numbers have positive Lebesgue measure. We expect that their union actually forms a set of full measure (see Figure (6)) and that this could be studied directly with (12) or using arguments developed in [14, 15].
@@
      $[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[0,\eta_{1}],$ (15)
@@
      $[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[\eta_{2},1],$ (16)
@@
The proposition ensures the following properties to the regions of parameters defined by (12). Suppose $\alpha\in[0,1-\rho)$. If $\rho\notin\mathbb{Q}$, the region contains exclusively values of $(\delta,a)$ corresponding to a map of the type (7). If $\rho\in\mathbb{Q}$ and $\alpha\neq 0$, the same remains true, excepted possibly for the value of $(\delta,a)$ given by the (right) extremity of each interval (12). This value of $(\delta,a)$ can belong to $M_{2}$ ($a\in[\eta_{1},\eta_{2}]$) instead of $M_{1}$ ($a<\eta_{1}$) and in such a case the corresponding map is of the type (9). If $\alpha=0$, the region is not necessarily contained in $M$, but its intersection with $M$ is nonempty and has the same properties as the regions with $\alpha\neq 0$. The same holds for $\alpha\in(1-\rho,1]$, replacing (7) by (8), $\alpha\neq 0$ by $\alpha\neq 1$, (right) by (left), $M_{1}$ ($a<\eta_{1}$) by $M_{3}$ ($a>\eta_{2}$) and $\alpha=0$ by $\alpha=1$. Finally, for $\alpha=1-\rho$ and $\rho\notin\mathbb{Q}$ the region contains exclusively values of $(\delta,a)$ corresponding to a map of the type (9). If $\rho\in\mathbb{Q}$, then the region still contains such values of $(\delta,a)$, but also intersects a large domain of $M_{1}$ and/or $M_{3}$. All these features can be observed in Figure 5.
Figure 5: Plot of the regions (12) for $\lambda=0.7$, $d=0.2$ and $\rho=1/5,1/4,1/3,2/5$, $1/2,3/5,2/3,3/4,4/5$ (from left in blue to right in red), with $\alpha\in[0,1-\rho)$ (left panel) $\alpha=1-\rho$ (middle panel) and $\alpha\in(1-\rho,1]$ (right panel).
@@
In the left panel of Figure 6, we plot the regions of parameters obtained with (12) for a set of $127$ values of $\rho$. If the set $M$ could be covered using a finite set of values of $\rho$, then some values of rotation number would not be reached in the family (3), which contradicts Theorem 2.3. This is why in Figure 6 we observe values of $(\delta,a)$ that do not belong to any of the plotted regions, see right panel. Therefore, the question we ask now is if for any $(\delta,a)\in M$, there exist some $\rho\in(0,1)$ and $\alpha\in[0,1]$ such that $\delta$ and $a$ satisfy (12). If the answer is in the positive, then by Theorem 2.3 and unicity of the rotation number, any map $f$ of the form (3) satisfies (12) for $\rho=\rho_{f}$ and some $\alpha\in[0,1]$. In other words, the reciprocal of Theorem 2.3 would be true.
@@
A direct corollary of Theorem 2.3 and its reciprocal Theorem 2.7, is that the rotation number of $f$ is $\rho\in(0,1)$ if and only if $\delta$ and $a$ satisfy (12) for some $\alpha\in[0,1]$.
@@
The monotonicity and continuity properties of $\phi_{\delta,\rho,\alpha}$ are given in Proposition 3.6. In particular, it is right continuous and non-decreasing. Besides, if $\rho\in(0,1)\setminus\mathbb{Q}$, then it is increasing and therefore admits a continuous generalized inverse.
@@
    with $\alpha\in[0,1-\rho]$ and $\eta=\eta_{1}$ if $(\delta,a)\in M_{1}$, $\alpha=1-\rho$ and $\eta=a$ if $(\delta,a)\in M_{2}$ and with $\alpha\in[0,1-\rho]$ and $\eta=\eta_{2}$ if $(\delta,a)\in M_{3}$.
@@
Theorem 2.7 ensures that any map $f=f_{\delta,a}$ of the type (7), (8) or (9) with no (ghost) fixed point satisfies (17). Therefore, if for instance $f_{\delta,a}$ is of the type $M_{1}$, then it satisfies (20) for $\rho=\rho_{f}$ and an $\alpha\in[0,1-\rho_{f}]$, unless $\rho_{f}\in\mathbb{Q}$ and $(\delta,a)$ is such that $\delta=\delta(\rho_{f},\alpha)$ or $a=a(\delta,\rho_{f}^{+},\alpha)$. We believe that the maps that do not satisfy (20) have actually a ghost periodic orbit and an empty attractor $\Lambda$, as it is the case for the contracted rotation (4) when $\rho_{f}\in\mathbb{Q}$ and $\delta=\delta(\rho_{f},1)$ (see, Theorem 7.2 in [16]). Let us suppose that $f_{\delta,a}$ satisfies the item 1) of Theorem 2.9 for some $\rho\in(0,1)$ and some $\alpha\in[0,1]$. Then, the map $f_{\delta,a}$ has rotation number $\rho$ (by Theorem 2.3) and the following additional properties:
@@
Now we give the interpretation of item 3) and of the parameter $\alpha$ with an example. Let us suppose that $(\delta,a)\in M_{1}$ (and therefore $\alpha\in[0,1-\rho]$ and $\eta=\eta_{1}(\delta)$). Then, consider the natural partition $\mathcal{P}_{a,\eta}$ of the phase space of $f_{\delta,a}$ defined by $I_{0}=[0,a)$, $I_{1}=[a,\eta)$ and $I_{2}=[\eta,1)$, and on the other hand the partition $\mathcal{P^{\prime}}_{\alpha,1-\rho}$ of the phase space of $R_{\rho}$ defined by $Y_{0}=[0,\alpha)$, $Y_{1}=[\alpha,1-\rho)$ and $Y_{2}=[1-\rho,1)$ (for $\alpha\in\{0,1-\rho\}$, the interval $Y_{0}$ or $Y_{1}$ is empty). Now, let $y\in[0,1)$ and define the code of the orbit of $y$ by $R_{\rho}$ with respect to $\mathcal{P^{\prime}}_{\alpha,1-\rho}$, as the sequence $i=\{i_{n}\}_{n\in\mathbb{N}}$ such that $R^{n}_{\rho}(y)\in Y_{i_{n}}$ for every $n\in\mathbb{N}$. Then, (22) states that $f_{\delta,a}^{n}(\phi_{\delta,\rho,\alpha}(y))\in I_{i_{n}}$ for every $n\in\mathbb{N}$, that is, the sequence $i$ is also the code of an orbit of $f_{\delta,a}$ with respect to $\mathcal{P}_{a,\eta}$. So, the set of the codes of $f_{\delta,a}$ restricted to the image of $\phi_{\delta,\rho,\alpha}$ is equal to the set of the codes of $R_{\rho}$ with respect to $\mathcal{P^{\prime}}_{\alpha,1-\rho}$. The same is also true for $(\delta,a)\in M_{2}$ but with $\eta=a$ and $\alpha=1-\rho$, as well as for $(\delta,a)\in M_{3}$ but with $\eta=\eta_{2}$ and the partitions $\mathcal{P}_{\eta,a}$ and $\mathcal{P^{\prime}}_{1-\rho,\alpha}$.
To sum up, Theorem 2.3 and Theorem 2.7 ensure that when $\alpha$ ranges over $[0,1]$, the intervals (20) provide all the maps with non-empty attractor of the form (3) with a given rotation number $\rho\in(0,1)$. The interpretation of the parameter $\alpha$ follows from item 3) of Theorem 2.9: the choice of $\alpha$ in the intervals (20) allows selecting, among all the maps with rotation number $\rho$, only maps whose set of codes contains those of the rotation $R_{\rho}$, with respect to the partition defined by $1-\rho$ and $\alpha$. In particular, by Proposition 2.5, the intervals (20) give maps of $M_{1}$ for the partitions such that $\alpha<1-\rho$, and of $M_{3}$ for those with $\alpha>1-\rho$. For $\alpha=1-\rho$, the maps belong to $M_{2}$ if $\rho\notin\mathbb{Q}$ and can belong to any models otherwise. Finally (23) in Theorem 2.9 states that the PCM with an attractor whose closure is a Cantor set are obtained by letting $\rho\notin\mathbb{Q}$ in (20), and those with $q$-periodic orbits by letting $\rho=p/q$. Then, the formula (19) of the function $\phi_{\delta,\rho,\alpha}$ allows obtaining the points of this attractor.
@@
In Figure 7, we plot in the left panel the graph of a rotation of angle $\rho=1/4$ and let $\alpha=5/16<1-\rho$ to obtain the partition $\mathcal{P^{\prime}}_{\alpha,1-\rho}=\{[0,5/16),[5/16,3/4),[3/4,1)\}$ of its phase space. Then, we let $\lambda=7/10$, $d=1/5$, choose $\delta=0.27$ and $a=0.34$ in the corresponding intervals (20) and plot $f_{\delta,a}$ in the right panel. We obtain a map of the type (7) ($a<\eta_{1}\simeq 0.757$), with rotation number $\rho=1/4$, an attractor composed of two periodic orbits of period 4, whose codes with respect to the partition $\{[0,a),[a,\eta_{1}),[\eta_{1},1)\}$ are those of the orbits of the rotation with respect to the partition $\mathcal{P^{\prime}}_{\alpha,1-\rho}$. The points of the two periodic orbits are given by the image of $\phi_{\delta,\rho,\alpha}$, see center panel.

### v3

@@ 2.2 Main results
Let $\lambda\in(0,1)$ and $d\in(0,1-\lambda)$. For any $\rho\in[0,1]$ and $\alpha\in[0,1]$, let $\delta(\rho,\alpha)$ and $a(\delta,\rho,\alpha)$ be defined by
@@
    We have $\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\subset(1-\lambda-d,1)$, and if $\delta$ satisfies (14), then $a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\cap[0,1]\neq\emptyset$ with $a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[0,1]$ if $\alpha\in(0,1)$ or $\rho\notin\mathbb{Q}$.
  2. 2.
@@
    if $k\rho\notin\mathbb{Z}+\alpha$ for every $k\in\mathbb{Z}^{*}$ and otherwise either $\delta(\rho^{-},\alpha)<\delta(\rho,\alpha)$ or $a(\delta,\rho^{+},\alpha)<a(\delta,\rho,\alpha)$.
  3. 3.
@@
Theorem 2.3 can be read as follows. Let $\rho\in(0,1)$. If $\alpha$ is any number in $[0,1]$ and $\delta$ is chosen in $\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]$, then, to obtain a map (6) with rotation number $\rho$ it is enough to select any $a\in[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\cap[0,1]$. Item 1) ensures that the region
$\mathcal{P}_{\rho,\alpha}:=\{(\delta,a)\in\mathbb{R}^{2}:\delta\in[\delta(\rho^{-},\alpha),\delta(\rho,\alpha)]\quad\text{and}\quad a\in[a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\}$ (17)
defined by (14) intersects our set of interest $M=(1-\lambda-d,1)\times[0,1]$ for every $\rho\in(0,1)$ and $\alpha\in[0,1]$, and is actually included in $M$ unless $\rho\in\mathbb{Q}$ and $\alpha\in\{0,1\}$.
@@
As shown by items 2) and 3), the size of the intervals (14) depends on the rational or irrational nature of $\rho$, but also on the relation between $\rho$ and $\alpha$. For $\rho$ irrational, the intervals (14) are both a singleton for almost all values of $\alpha$. These intervals are not singletons only if $\alpha$ is such that $\{k\rho\}=\alpha$ for some $k=k_{\alpha}\in\mathbb{Z}^{*}$, that is, if $\alpha$ belongs to the orbit of $\rho$ by the irrational rotation of the same angle. For such $\alpha$, one of the two intervals (but not both) is not a singleton. Its exact length depends on $k_{\alpha}$ and is given in Lemma 4.7. For $\rho=p/q$ rational, the intervals (14) have always positive Lebesgue measure. The inequalities (15) and (16) become equalities if $\alpha$ does not belong to the $q$-periodic orbit of $\rho$ by the rotation of the same angle. In particular, if $p$ and $q$ are coprime, the size of both intervals is a function of $\alpha$ which is constant in $0,1]\setminus\{0,1/q,2/q,\dots,1\}$. Now, for $\alpha\in\{0,1/q,2/q,\dots,1\}$ the intervals are bigger and their exact length, which depends on $\alpha$, is given in Lemma 4.8.
If $\rho=p/q$ is rational, the series defining $\delta(\rho,\alpha)$ and $a(\delta,\rho,\alpha)$ are finite sums given in Lemma 4.9, which can be numerically computed with a high precision. Also, $\delta(\rho^{-},\alpha)$ and $a(\delta,\rho^{+},\alpha)$ are deduced from item 3), or from Lemma 4.8 if $rp/q\in\mathbb{Z}+\alpha$ for some $r\in\{1,\dots,q-1\}$. As an example, we carry out the numerical computation of $\mathcal{P}_{\rho,\alpha}$ for $\rho=1/3$ when $\alpha=1/2$ and plot it in the left panel of Figure 4. Since $a(\delta,\rho,\alpha)$ is an affine function of $\delta$ and $a(\delta,\rho,\alpha)-a(\delta,\rho^{+},\alpha)$ does not depend on $\delta$ (see item 3) or Lemma 4.8), the region obtained is a parallelogram. Now, as $\delta(p/q,\alpha)$ and $a(\delta,p/q,\alpha)$ are piecewise constant functions of $\alpha$, see Lemma 4.9, if $\alpha$ takes all the values in $[0,1]$, then (14) gives only a finite number of different regions $\mathcal{P}_{\rho,\alpha}$ for which the rotation number is $p/q$. If $p$ and $q$ are coprime, this number is equal to $2q+1$ and the regions are obtained by letting $\alpha=l/q$ for every $l\in\{0,1,\dots,q\}$ and by choosing any $\alpha\in(l/q,(l+1)/q)$ for every $l\in\{0,1,\dots,q-1\}$. In the right panel of Figure 4, the $2q+1=7$ regions $\mathcal{P}_{\rho,\alpha}$ corresponding to $\rho=1/3$ are shown. We note that $\mathcal{P}_{\rho,l/q}$ and $\mathcal{P}_{\rho,(l+1)/q}$ have a non-empty intersection which is equal to the region $\mathcal{P}_{\rho,\alpha}$ obtained for any $\alpha\in(l/q,(l+1)/q)$.
Figure 4: In green, the set $\mathcal{P}_{\rho,\alpha}$ for $\lambda=0.7,d=0.2$, $\rho=p/q=1/3$ and different values of $\alpha$. Left panel: the parallelogram $\mathcal{P}_{1/3,1/2}$. Right panel: the union of $\mathcal{P}_{1/3,\alpha}$ over all $\alpha\in[0,1]$. The resulting set is made of $2q+1=7$ different regions $\mathcal{P}_{1/3,\alpha}$. We note that $\mathcal{P}_{1/3,1/2}$ is also given by the intersection $\mathcal{P}_{1/3,1/3}\cap\mathcal{P}_{1/3,2/3}$.
@@
      $a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[0,\eta_{1}],$ (18)
@@
      $a(\delta,\rho^{+},\alpha),a(\delta,\rho,\alpha)]\subset[\eta_{2},1],$ (19)
@@
The proposition ensures the following properties to the regions $\mathcal{P}_{\rho,\alpha}$. Suppose $\alpha\in[0,1-\rho)$. If $\rho\notin\mathbb{Q}$, the region $\mathcal{P}_{\rho,\alpha}$ contains exclusively values of $(\delta,a)$ corresponding to a map of the type (9). If $\rho\in\mathbb{Q}$ and $\alpha\neq 0$, the same remains true, excepted possibly for the value of $(\delta,a)$ given by the (right) extremity of each interval (14). This value of $(\delta,a)$ can belong to $M_{2}$ ($a\in[\eta_{1},\eta_{2}]$) instead of $M_{1}$ ($a<\eta_{1}$) and in such a case the corresponding map is of the type (11). If $\alpha=0$, the region $\mathcal{P}_{\rho,\alpha}$ is not necessarily contained in $M$, but its intersection with $M$ is nonempty and has the same properties as the regions with $\alpha\neq 0$. The same holds for $\alpha\in(1-\rho,1]$, replacing (9) by (10), $\alpha\neq 0$ by $\alpha\neq 1$, (right) by (left), $M_{1}$ ($a<\eta_{1}$) by $M_{3}$ ($a>\eta_{2}$) and $\alpha=0$ by $\alpha=1$. Finally, for $\alpha=1-\rho$ and $\rho\notin\mathbb{Q}$ the region $\mathcal{P}_{\rho,\alpha}$ contains exclusively values of $(\delta,a)$ corresponding to a map of the type (11). If $\rho\in\mathbb{Q}$, then $\mathcal{P}_{\rho,\alpha}$ still contains such values of $(\delta,a)$, but also intersects a large domain of $M_{1}$ and/or $M_{3}$. All these features can be observed in Figure 5.
Figure 5: Plot of the $\mathcal{P}_{\rho,\alpha}$ for $\lambda=0.7$, $d=0.2$ and $\rho=1/5,1/4,1/3,2/5$, $1/2,3/5,2/3,3/4,4/5$ (from left in blue to right in red), with $\alpha\in[0,1-\rho)$ (left panel) $\alpha=1-\rho$ (middle panel) and $\alpha\in(1-\rho,1]$ (right panel).
@@
In the left panel of Figure 6, we plot the union over $\alpha\in[0,1]$ of $\mathcal{P}_{\rho,\alpha}$ for a set of $127$ values of $\rho$. If the set $M$ could be covered using a finite set of values of $\rho$, then some values of rotation number would not be reached in the family (6), which contradicts Theorem 2.3. This is why in Figure 6 we observe values of $(\delta,a)$ that do not belong to any of the plotted regions, see right panel. Therefore, the question we ask now is if for any $(\delta,a)\in M$, there exist some $\rho\in(0,1)$ and $\alpha\in[0,1]$ such that $\delta$ and $a$ satisfy (14). If the answer is in the positive, then by Theorem 2.3 and unicity of the rotation number, any map $f$ of the form (6) satisfies (14) for $\rho=\rho_{f}$ and some $\alpha\in[0,1]$. In other words, the reciprocal of Theorem 2.3 would be true.
@@
A corollary of Theorem 2.3 and its reciprocal Theorem 2.6 is the following.
Corollary 2.7.
Let $\lambda\in(0,1)$ and $d\in(1-\lambda,1)$. For any $\delta\in(1-\lambda-d,1)$ and $a\in[0,1]$ let $f_{\delta,a}$ be the map (6).
  1. 1.
    For any $\rho\in(0,1)$, the rotation number of $f_{\delta,a}$ is $\rho$ if and only if $(\delta,a)\in\bigcup_{\alpha\in[0,1]}\mathcal{P}_{\rho,\alpha}$.
  2. 2.
    The set $\mathcal{I}:=\{(\delta,a)\in(1-\lambda-d,1)\times[0,1]:\text{the rotation number of }f_{\delta,a}\text{ is irrational}\,\}$ is a Lebesgue measurable set.
Proof.
Item 1) follows directly from (20) and Theorem 2.3. Now, let $\mathcal{P}_{0,\alpha}:=\mathcal{F}_{1}\cup\mathcal{F}_{2}$ for every $\alpha\in[0,1]$. Then, from item 1) we deduce that
$M\setminus\mathcal{I}=\bigcup_{\rho\in\mathbb{Q}\cap[0,1)}\bigcup_{\alpha\in[0,1]}(\mathcal{P}_{\rho,\alpha}\cap M).$
For any $\rho\in\mathbb{Q}\cap[0,1)$ and $\alpha\in[0,1]$ the set $\mathcal{P}_{\rho,\alpha}$ is Lebesgue measurable. On the other hand, for any $\rho\in\mathbb{Q}\cap[0,1)$ the set $\{\mathcal{P}_{\rho,\alpha}:\alpha\in[0,1]\}$ is finite, since $\delta(\rho,\alpha)$ and $a(\delta,\rho,\alpha)$ are piecewise constant functions of $\alpha$, see Lemma 4.9. It follows that $\mathcal{I}$ is the complement in $M$ of a countable union of Lebesgue measurable sets, which proves item 2). ∎
Remark 2.8.
Item 2) of Theorem 2.3 shows that for $\rho$ irrational and $\alpha$ in the orbit of $\rho$ by the rotation of angle $\rho$, the region $\mathcal{P_{\rho,\alpha}}$ is a segment. It follows that the set $\mathcal{I}$ of the values of $(\delta,a)$ for which the rotation number of $f_{\delta,a}$ is irrational has Hausdorff dimension at least equal to 1. On the other hand, item 2) of Corollary 2.7 together with the results of [24] allow to prove that this set has zero Lebesgue measure, as shown in the following. Since we have proved that $\mathcal{I}$ is a Lebesgue measurable set, we can apply Fubini’s theorem to obtain that
$\ell_{2}(\mathcal{I})=\int_{\mathbb{R}^{2}}\mathbb{1}_{\mathcal{I}}(\delta,a)\ d\ell_{2}(\delta,a)=\int_{\mathbb{R}}\left(\int_{\mathbb{R}}\mathbb{1}_{\mathcal{I}_{a}}(\delta)\ d\ell_{1}(\delta)\right)\ d\ell_{1}(a),$ (21)
where $\ell_{n}$ denotes the Lebesgue measure in $\mathbb{R}^{n}$ and
$\mathcal{I}_{a}:=\{\delta\in(1-\lambda-d,1):(\delta,a)\in\mathcal{I}\}\qquad\forall a\in[0,1].$
According to Theorem 1.1 of [24], for any $a\in[0,1]$ fixed, the set of the values of $\delta\in(1-\lambda-d,1)$ such that $f_{\delta,a}$ is not asymptotically periodic has zero Lebesgue measure. As $f_{\delta,a}$ has no periodic orbit if its rotation number is irrational, Theorem 1.1 of [24] implies that $\ell_{1}(\mathcal{I}_{a})=0$ for any $a\in[0,1]$ and we deduce from (21) that $\ell_{2}(\mathcal{I})=0$.
@@
The monotonicity and continuity properties of $\phi_{\delta,\rho,\alpha}$ are given in Proposition 3.6. It is right continuous and non-decreasing. Besides, if $\rho\in(0,1)\setminus\mathbb{Q}$, then it is increasing and admits a continuous generalized inverse.
@@
    with $\alpha\in[0,1-\rho]$ and $\eta=\eta_{1}$ if $(\delta,a)\in M_{1}$, $\alpha=1-\rho$ and $\eta=a$ if $(\delta,a)\in M_{2}$ and with $\alpha\in[1-\rho,1]$ and $\eta=\eta_{2}$ if $(\delta,a)\in M_{3}$.
@@
Theorem 2.6 ensures that any map $f=f_{\delta,a}$ of the type (9), (10) or (11) with no (ghost) fixed point satisfies (20). Therefore, if for instance $f_{\delta,a}$ is of the type $M_{1}$, then it satisfies (24) for $\rho=\rho_{f}$ and an $\alpha\in[0,1-\rho_{f}]$, unless $\rho_{f}\in\mathbb{Q}$ and $(\delta,a)$ is such that $\delta=\delta(\rho_{f},\alpha)$ or $a=a(\delta,\rho_{f}^{+},\alpha)$. We believe that the maps such that $\rho_{f}\in\mathbb{Q}$ and $\delta=\delta(\rho_{f},\alpha)$ or $a=a(\delta,\rho_{f}^{+},\alpha)$ have a ghost periodic orbit and an empty attractor $\Lambda$, as it is the case for the contracted rotation (1) when $\rho_{f}\in\mathbb{Q}$ and $\delta=\delta(\rho_{f},1)$ (see Theorem 7.2 in [19]). Let us suppose that $f_{\delta,a}$ satisfies the item 1) of Theorem 2.10 for some $\rho\in(0,1)$ and some $\alpha\in[0,1]$. Then, the map $f_{\delta,a}$ has rotation number $\rho$ (by Theorem 2.3) and the following additional properties:
@@
Now we give the interpretation of item 3) and of the parameter $\alpha$ with an example. Let us suppose that $(\delta,a)\in M_{1}$ (and therefore $\alpha\in[0,1-\rho]$ and $\eta=\eta_{1}(\delta)$). Then, consider the natural partition $P_{a,\eta}$ of the phase space of $f_{\delta,a}$ defined by $I_{0}=[0,a)$, $I_{1}=[a,\eta)$ and $I_{2}=[\eta,1)$, and on the other hand the partition $P^{\prime}_{\alpha,1-\rho}$ of the phase space of $R_{\rho}$ defined by $Y_{0}=[0,\alpha)$, $Y_{1}=[\alpha,1-\rho)$ and $Y_{2}=[1-\rho,1)$ (for $\alpha\in\{0,1-\rho\}$, the interval $Y_{0}$ or $Y_{1}$ is empty). Now, let $y\in[0,1)$ and define the code of the orbit of $y$ by $R_{\rho}$ with respect to $P^{\prime}_{\alpha,1-\rho}$, as the sequence $i=\{i_{n}\}_{n\in\mathbb{N}}$ such that $R^{n}_{\rho}(y)\in Y_{i_{n}}$ for every $n\in\mathbb{N}$. Then, (26) states that $f_{\delta,a}^{n}(\phi_{\delta,\rho,\alpha}(y))\in I_{i_{n}}$ for every $n\in\mathbb{N}$, that is, the sequence $i$ is also the code of an orbit of $f_{\delta,a}$ with respect to $P_{a,\eta}$. So, the set of the codes of $f_{\delta,a}$ restricted to the image of $\phi_{\delta,\rho,\alpha}$ is equal to the set of the codes of $R_{\rho}$ with respect to $P^{\prime}_{\alpha,1-\rho}$. The same is also true for $(\delta,a)\in M_{2}$ but with $\eta=a$ and $\alpha=1-\rho$, as well as for $(\delta,a)\in M_{3}$ but with $\eta=\eta_{2}$ and the partitions $P_{\eta,a}$ and $P^{\prime}_{1-\rho,\alpha}$.
Remark 2.11.
As shown in the right panel of Figure 4, for $\rho=p/q$ rational the values of $\delta$ and $a$ satisfying (24) for $\alpha\in(l/q,(l+1)/q)$, where $l\in\{0,\dots,q-1\}$, also satisfy (24) for $\alpha^{\prime}=l/q$. It follows that (27) holds with $\phi_{\delta,\rho,\alpha}$ and $\phi_{\delta,\rho,\alpha^{\prime}}$. As $\phi_{\delta,\rho,\alpha}([0,1))$ is the union of two periodic orbits and $\phi_{\delta,\rho,\alpha^{\prime}}([0,1))$ is a unique periodic orbit, we deduce that the inclusion (27) can indeed be strict.
@@
As a first application of Theorem 2.10, we show how to construct a map of the type (9) with 2 periodic orbits of period $4$. Choose any $\lambda\in(0,1)$, $d\in(0,1-\lambda)$ and let $\rho=1/4$. Now, let $\alpha=5/16<1-\rho$ and choose $\delta$ and $a$ in the corresponding intervals (24). Then $f_{\delta,a}$ is of the type (9), has rotation number $\rho=1/4$, an attractor composed of two periodic orbits, since $\alpha\not\in\{0,1/4,1/2,3/4,1\}$, whose codes with respect to the partition $\{[0,a),[a,\eta_{1}),[\eta_{1},1)\}$ are those of the orbits of the rotation with respect to the partition $\{[0,\alpha),[\alpha,1-\rho),[1-\rho,1)\}$. A numerical example is given in Figure 7.